OpenStudy (anonymous):

Which ordered pair is NOT a solution of the inequality, 3x - 2y ≤ 12 A.(-3, -10.5) B.(-1, -7) C.(0, -6) D.(2, -3)

OpenStudy (anonymous):

none..they are all solutions. subbing in the first gives -9 + 21 = 12 which is fine, subbing in the second gives -3 + 13 which is 11 < 12 which is fine, the third: 12=12 which is fine and the fourth: 6+6=12 which is also fine.

OpenStudy (anonymous):

Thats what i said but thats not an option :/

OpenStudy (anonymous):

I meant on the second one: -3 + 14 = 11* still doesn't change it though